OpenStudy (jocelynstalnaker):

4 question project. Help?

OpenStudy (jocelynstalnaker):

Iris has been studying an invasive population of snails. This particular snail has no local predators, so the population grows wildly. She has observed that the population follows an exponential rate of growth for fifteen years. 1. Create your own exponential function, f(x), which models the snail population. You will need to identify the principal population of the snails and the rate of growth each year. Explain to Iris how your function shows the principal population and the rate of growth, in complete sentences. 2. A local snail population grows according to the function g(x) = 200(1.03)2x. Demonstrate the steps to convert g(x) into an equivalent function with only x as the exponent. Then, explain to Iris how the key features of this local snail population compares to the key features of the invasive population. 3. Iris wants to graph the invasive snail population to show the city council. Justify what the appropriate domain and range would be for the function f(x), what the y-intercept would be, and if the function is increasing or decreasing. 4. In five years, a garden festival plans on using the park where Iris has been studying the invasive snails. Explain to the garden festival committee how to find the average rate of change for the snail population between years 2 and 5. Describe what this average rate of change represents.

zepdrix (zepdrix):

Ahhh so many wordsssss :O(

zepdrix (zepdrix):

Number 1 is pretty straight forward: `create your own`, do you know how to do that?

OpenStudy (jocelynstalnaker):

I know >_>

OpenStudy (jocelynstalnaker):

This is what i put (1) The exponential equation with the population starting at 45 and doubling each year would be f(x)=45(2)^x where x is the year

zepdrix (zepdrix):

Ooo very nice!

OpenStudy (jocelynstalnaker):

ye

zepdrix (zepdrix):

So where we stuck? Number 2?

OpenStudy (jocelynstalnaker):

yes

zepdrix (zepdrix):

\[\Large\rm g(x) = 200(1.03)^{2x}\]Hmm I guess they want us to use some exponent rules...

zepdrix (zepdrix):

One of our exponent rules allows us to do this, yes? \[\Large\rm g(x) = 200\left[(1.03)^{x}\right]^2\]

zepdrix (zepdrix):

I'm not sure if that's the one we want to use though, hmm...

zepdrix (zepdrix):

Our other option would be to use the addition rule.. thing...\[\Large\rm g(x) = 200(1.03)^{x+x}=200(1.03)^x\cdot(1.03)^x\]But then again, that just leads us to the square.

zepdrix (zepdrix):

So what would we say has changed about the key features of g(x)? +_+ Hmmm

zepdrix (zepdrix):

Something like.... how the growth rate is being squared? So the population will increase much faster.

zepdrix (zepdrix):

It doesn't change the initial population or anything like that.

zepdrix (zepdrix):

Do you understand the exponent rule I used? :U

OpenStudy (jocelynstalnaker):

Sorry, I had to eat! @zepdrix

OpenStudy (jocelynstalnaker):

I dont really understand any of this v_v

zepdrix (zepdrix):

D:

OpenStudy (jocelynstalnaker):

I know, I'm sorry ;(

zepdrix (zepdrix):

Exponent rule tells us:\[\LARGE\rm (a^{\color{orangered}{b}})^{\color{royalblue}{c}}=a^{\color{orangered}{b}\color{royalblue}{c}}\]We just multiply the exponents. We want to use this rule in reverse,\[\LARGE\rm g(x) = 200(1.03)^{\color{orangered}{2}\color{royalblue}{x}}\]\[\LARGE\rm g(x) = 200\left[(1.03)^{\color{royalblue}{x}}\right]^{\color{orangered}{2}}\]

OpenStudy (jocelynstalnaker):

Ok

zepdrix (zepdrix):

So with our function:\[\Large\rm g(x) = 200\left[(1.03)^{x}\right]^2\]This is the population:\[\Large\rm g(x) = \color{orangered}{200}\left[(1.03)^{x}\right]^2\]And this is what's making the population grow:\[\Large\rm g(x) = 200\color{orangered}{\left[(1.03)^{x}\right]^2}\]Soooo, for problem 2, maybe say something likeeeeee... The growth rate is being squared, which causes it to grow much faster. I dunno something like that :c

zepdrix (zepdrix):

For problem 3: The y-intercept is the `initial population`. Notice that at the very start Iris said that the growth rate of these snails follows an exponential curve for a period of `15 years`. That tells us our domain, our x values that we should use. A good domain would be something like \(\Large\rm 0\le x\le 15\), yes?

zepdrix (zepdrix):

The snails would `range` from the initial population up to whatever value they reach at 15 years.

OpenStudy (jocelynstalnaker):

Yes, Thank you so much!

OpenStudy (jocelynstalnaker):

And so number 4 I would just replace x with 5 in my equation?

zepdrix (zepdrix):

Remember how to find `average rate of change` or `slope` ?\[\Large\rm m=\frac{y_2-y_1}{x_2-x_1}\]Our y_2 is the 5 year mark, while our y_1 is the 2 year mark. I'll rewrite the slope formula using function notation, because that's what's going to help us here:\[\Large\rm m=\frac{g(x_2)-g(x_1)}{x_2-x_1}\]Plugging in our 5 and 2 for x,\[\Large\rm m=\frac{g(5)-g(2)}{5-2}\]

zepdrix (zepdrix):

So before you can calculate your slope/rate of change, you need to figure out the population at 5 years, at 2 years, and plug those in.

OpenStudy (jocelynstalnaker):

So m=g(3)/3

zepdrix (zepdrix):

Our y_2 is the `population at the` 5 year mark, blah i said that a little sloppy the first time.

zepdrix (zepdrix):

No no no no no.

OpenStudy (jocelynstalnaker):

Or am I doing this wrong

zepdrix (zepdrix):

\[\Large\rm g(5)-g(2)\ne g(3)\]I'm not sure what gave you the idea that you could do that :c

OpenStudy (jocelynstalnaker):

Ahh Sorry

OpenStudy (jocelynstalnaker):

Wait, thats what I put >_> I just put it over 3

zepdrix (zepdrix):

Yes, do you see how I put a `not equal to` symbol between them? :o That is not correct to do. The 3 in the bottom is fine though. \[\LARGE\rm g(\color{royalblue}{x}) = 200(1.03)^{2\cdot\color{royalblue}{x}}\]\[\LARGE\rm g(\color{royalblue}{5}) = 200(1.03)^{2\cdot\color{royalblue}{5}}\]

zepdrix (zepdrix):

Here is how you would get your g(5) value. Calculator should help out.

zepdrix (zepdrix):

So what do you get for g(5)? +_+

OpenStudy (jocelynstalnaker):

2060

OpenStudy (jocelynstalnaker):

Right?

zepdrix (zepdrix):

No not even close :c Hmm let's make sure we're putting it into the calculator correctly.

zepdrix (zepdrix):

\[\Large\rm 200(1.03)\text{^}(10)\]Put it in like that.

OpenStudy (jocelynstalnaker):

Yes, I did

OpenStudy (jocelynstalnaker):

OH WAIT

OpenStudy (jocelynstalnaker):

I had a blonde moment.. even though I'm not a blonde.. Wow, ok.. Exponent, duhh

zepdrix (zepdrix):

lol :p

OpenStudy (jocelynstalnaker):

268.78

zepdrix (zepdrix):

Mmm ya there we go! c:\[\Large\rm m=\frac{268.78-g(2)}{5-2}\]

zepdrix (zepdrix):

How bout g(2)? Think you can figure that one out? c:

OpenStudy (jocelynstalnaker):

268.78-\[268.78 - 1.25\div5-2\]1.125

OpenStudy (jocelynstalnaker):

1.125 >_<

OpenStudy (jocelynstalnaker):

I cant do the equation thing.. lol

zepdrix (zepdrix):

Your g(2) looks .... kinda crazy. If g(0) represents the initial population.... 200 snails and g(5) represents the snail population after 5 years .... 268 snails do you think a population of 1 is about right after 2 years?

zepdrix (zepdrix):

After 2 years, we would expect the population to be a little more than 200, since it's increasing. So I dunno how you got that number >.<

zepdrix (zepdrix):

\[\Large\rm g(2)=200(1.03)\text{^}(4)\]

OpenStudy (jocelynstalnaker):

I forgot to multiply the 200... 225.101

zepdrix (zepdrix):

Oh ok.\[\Large\rm m=\frac{268.78-225.10}{5-2}\]Great, so wrap it up with one more calculation!

OpenStudy (jocelynstalnaker):

m=14.56

zepdrix (zepdrix):

Yay good job! That's telling us that over a 3 year period, the snail population was increasing at a rate of 14 and a half snails per year.